C = 2350 * 10^(0.099 v). When C=3750, find v. can you help me
Answer by Cromlix(4381)

C = 2350 * 10^(0.099 v)
When C = 3750, find v
C = 2350 * 10^(0.099 v)
3750 = 2350*10^(0.099v)
3750/2350 = 10^(0.099v)
1.5957 = 10^(0.099v)
log10 1.5957 = 0.099v log10 10
(log10 10 = 1)
log10 1.5957/0.099 = v
v = 2.05
Hope this helps.
:-)